45* . Prove that if a, b, c are three different integers, then there exist in-

finitely many positive integers n such that a+n, b+n, c+n are pairwise rel-

atively prime. 


1
Expert's answer
2022-06-20T02:42:31-0400

If the integers "a,b,c" are distinct, then the number "h=(a-b)(a-c)(b-c)" is different from zero. In case "h\ne\pm1", let "q_1,\dots,q_n" denote all prime ">3" divisors of "h".


If two or more among numbers "a,b,c" are even, put "r=1", otherwise put "r=0". Clearly, at least two of the numbers "a+r", "b+r", "c+r" will be odd. If "a,b,c" give three diferent remainders upon dividing by "3", put "r_0". If two or more among "a,b,c" give the same remainder "\rho" upon dividing by "3", put "r_0=1-\rho". Clearly, at least two of the number "a+r_0, b+r_0, c+r_0" will be not divisible by "3".


Now, let "i" denote one of the numbers "1,2,\dots, s". Exists an integer "r_i" such that none of the numbers "a+r_i, b+r_i, c+r_i" is divisible by "q_i". According to the Chinese remainder theorem, there exist infinitely many positive integer "n" such that "n\equiv r\mod2", "n=r_0\mod3", and "n\equiv r_i\mod q_i" for "i=1,2,\dots, s".


We shall show that the numbers "a+n,b+n" and "c+n" are pairwise relatively prime. Suppose, for instance, that "(a+n_,b+n)>1". Then  there would exist a prime "q" such that "q|a+n" and "q|b+n", hence "q|a-b", which implies "q|h" and "h\ne\pm1". Since "n\equiv r\mod 2" and at least two of the numbers "a+n, b+n, c+n" are odd, and we cannot have "q=2". Next, since "n\equiv r_0\mod3" at least two of the numbers "a+r_0,b+r_0,c+r_0" are not divisible by 3, at least two of the numbers "a+n,b+n,c+n" are not divisible by "3", and we cannot have "q=3".


Since "q|h", in view of the definition of "h", we have "q=q_i" for a certain "i" from the sequence "1,2,\dots, s". However, in view of "n\equiv r_i\mod q_i", or "n\equiv r_i\mod q", and in view of the fact that none of the numbers "a+n,b+n,c+n" is divisible by "q_i=q", contrary to the assumption that "q|a+n" and "q|b+n". Thus, we provet that "(a+n,b+n)=1". In a similat way we show that "(a+n,c+n)=1", and "(b+n,c+n)=1". Therefore the numbers "a+n,b+n" and "c+n" are pairwise relatively prime. Since there are infinitely many such numbers "n", the proof is complete.